xx

Interlocking System Guide

 00

1 Introduction

In real-world railways, a so-called interlocking system is a set of railway signals and trackside equipment. Its purpose is to prevent conflicting train movements which otherwise could result in derailing or colliding trains. If you want more information, just search for “railway interlocking” on the internet. Real-world interlocking systems perform this task by setting routes. A route is a path along a track that a train can safely pass. To set a route for a train, the signalman (the operator of a signal box) has to set switches (turnouts) to the correct position and lock them in order to make a signal for a train show “Proceed”. In newer systems, this is done automatically by the interlocking system. A route can not be set if switches are locked to a wrong position by another route or if any portion of the route is occupied by a train. The interlocking system in this Minetest mod tries to follow real-world interlocking systems as far as applicable. It divides tracks into track sections and implements a route setting mechanism following the same principle. However, for the sake of simplicity of implementation and usage, not all concepts of real-world interlocking have been taken over. Especially, there is no mechanism for overlap. If you are looking for a place to learn how real-world interlocking systems work, have a look at “SimSig”. By looking at their simulations, you can obtain experience on how to set up your own interlocking systems in AdvTrains. The SimSig glossary is a good place to look up unknown terms in this document.

2 Setting

up track sections

In the real world, a line of track is divided into so-called track sections, or track circuits. Those systems often can not tell where exactly a train is, but only which track sections it occupies. A route can never be set through an occupied track section. A track section often covers:

A section on a main running line, between two signals

A single turnout

A rail crossing, or a set of turnouts acting as a double/single slip switch

A siding

You will find some examples on how to interlock certain patterns later.

2.1 Track Circuit Breaks

In this mod, you will not directly configure the locations of track sections. Instead, you designate the borders of each track section using a special node, the Track Circuit Break, abbreviated TCB. For example, if you want to create a track section for a piece of a main running line, you set up two TCBs at the ends of this track circuit. Setting up a TCB works as follows:

Place a TCB node somewhere near the place where the circuit break is going to be located.

Right-click the TCB node

Punch the rail which should act as TCB

The result should look like this: PIC Now you have assigned the TCB node to a rail. Right-click the TCB node once again. This will bring up a form which looks as follows: PIC You see that the form is divided in side A and side B. To designate where each side is, a marker is displayed on the rail. You can always make this marker show up by punching the TCB node, and remove it by punching the marker. Both sides are shown as “End of interlocking”. This means that there is no track section set up at this place. You should repeat this procedure once again a few meters away from the first TCB to create a second TCB on the same track. PIC Once you have both bordering TCBs set up, you can now create the actual track section. To do this:

Right-click one of the TCBs

Locate the correct side (A or B) to create the track section

Click “Create interlocked Track Section” in the formspec on the chosen side.

Now, the text on the formspec has changed. It shows something like this: PIC Clicking “Show Track Section” brings up another formspec: PIC On the top, you see a list of all TCBs that border this track section. In your case, there should be two TCBs listed. If there’s only one, head over to 2.2. You should now select a name for the track section, to identify it later. The same procedure is applicable when you create a turnout track section, except that you have to set up three or more TCBs. The AdvTrains interlocking system allows you to add more TCBs after you have created a track section. This works without problems in most cases. For example, you can easily insert a turnout into an already set-up track section and create another TCB behind it, and AdvTrains will automatically detect the existing track section. Problems arise only if you try to insert a TCB in-between a section, in which case both sides of the TCB will end up assigned to the same section. The code currently does not handle this case properly, so try to avoid this situation by all means. As a last resort, you can always dissolve a faulty track section, as described in the next chapter.

2.2 Long track sections, crossings and other edge cases

2.2.1 Very long track sections

If you try to set up a track section that is longer than 1000 nodes, advtrains won’t recognize the TCB at the other end because of a safety limit in the traverser function, which is supposed to prevent deadlocks. This case has happened when the Track Section overview screen only shows one TCB in the list. The procedure for this is as follows:

Go to the second TCB (the one that wasn’t recognized). It should show “End of Interlocking” on the relevant side.

Click “Create interlocked track section”. The section created will be different from the one that is already present.

In the track section overview, click “Join into other section”

Go back to the first TCB, bring up the Track Section overview screen of the first track section and click “Join with ???”

The other, missing TCB should now appear in the list. If you accidentally started such a joining procedure, click the “X” button on the right.

2.2.2 Rail crosses

Since rail crosses are created by laying tracks across each other without logical connection, there’s no way for advtrains to know whether rails cross each other. Rail crossings in interlocking systems are always one single track section, which in most cases has 4 TCBs adjacent. PIC The procedure is quite similar to the one for long sections: First, create two track sections for the branches, and then use the “Join” function to merge both sections into one.

2.2.3 Deleting and re-adding single TCBs to a section

In some occasions, for example when you remove a siding or a crossover, it can be necessary to unassign a TCB from a track section. There are multiple ways to do this:

In the TCB form, click the “Remove from section” button

In the track section form, first select the TCB in the list and then click “Unlink selected TCB”

The result is that the TCB shows “End of Interlocking” and the section does not list the TCB as an endpoint anymore. The other case is adding a siding or a crossover, in which case one or more TCBs still show “End of Interlocking” although they should be part of a section:

Go to another TCB that is registered in the track section and click “Update near TCBs”

If that did not work, follow the procedure of creating a long track section

2.2.4 Dissolving sections

If you made a mistake setting up something and you don’t see any other way to fix a misconfigured track section, you can always delete it using the “Dissolve section” button. This operation removes the track section and sets all TCBs that previously belonged to the section as “End of Interlocking”. This will always work and lets you start over new with setting up track sections.

2.3 Interlocking patterns

This section is supposed to show some examples on how you should set up track sections on certain track configurations.

3 Signals and routes

Signals are appliances that can give instructions to trains. That can be the permission to proceed, a speed restriction, or other information. There are 2 types of signals:

Static signals always display the same information to the train. This can be a speed restriction (or the end of one), a disallowal to proceed as shunt move or similar things. In most cases, these are signs.

Variable signals are what most people would call a “signal”. Its function is to inform trains about whether and at which speed they can proceed into the next section safely.

3.1 Signal Influence Point

Every signal is associated to a track on which the instruction should be followed. Signals are usually placed right next to the track on the right side. Human observers do know then that the signal belongs to the track left of it, however, train safety systems (like the one in advtrains) can not. This is the reason why a so-called “influence point” needs to be assigned to any signal that should actually give instructions to trains, should the driver (if even there is one) fail to recognize the instructions. Depending on the signal and the mod that adds the signal, there are different ways to configure this. Signals integrated into advtrains behave as follows:

Static signals and all red-green light signals from core advtrains that are not assigned to a TCB can be configured by holding the “Sneak” key and then right-clicking the signal

All signals that are assigned to a TCB can be configured by first right-clicking them, then selecting “Influence Point” in the signalling formspec.

The small formspec that opens allows you to set and later view or clear the Influence Point. To set the influence point, click the “Set” button, face towards the signal and punch a rail about 2m in front of the signal. A small marker will be shown, indicating success. To cancel setting an influence point, punch anything other. (note that then the influence point remains unset, regardless of its previous state) The advtrains-internal train safety system ensures that the train always obeys any restrictions imposed by signals, if (and only if) the influence point is set properly.

3.2 Main and Shunt signals

While static signals are mainly used for speed restrictions, the interesting ones are variable signals. Of course, you can always control any variable signal by traditional means (mesecons, digiline, right-click) if the signal allows it, but that misses the point of this interlocking system. In the following sections, we will talk about main signals. By this, we mean a variable signal that can display both a “Danger” aspect (trains are not allowed to proceed) and at least one “Proceed” aspect (train may proceed as train/shunt move, with optional speed restriction), which act as an “entry signal” for one or multiple routes.

3.3 The concept of routes

A so-called route is a locked path between two main signals, which locks all turnouts in the correct position. Its purpose is to offer a train a path on which it can safely proceed without interfering with any other train. A route always incorporates and locks one to multiple track sections, starting with the one that lies directly behind the “entry” signal.

Example: Imagine a station with 2 platforms on a single track running line. We are looking at signal A. You probably want trains coming from the right to go into platform 1 or into platform 2, so you need to program 2 routes. PIC This leads us to the most important aspect of route programming: Routes always start at a signal (A) and end at a signal facing in the same direction (D and E), not at an opposite-facing signal (B and C). There are only few exceptions, we’ll cover this later. When you set a route to make a train proceed on it, the interlocking system ensures that:

There are no rail vehicles on the route

All turnouts are set to the correct position and it is impossible to move them

No other routes can be set that would in any way conflict with this route

For this to work, you need to specify all track sections the train will pass along, as well as the positions of all turnouts that need to be locked. Those are not only the turnouts that lay directly on the train’s route, but also some turnouts on adjacent tracks, the so-called flank protection. The purpose of flank protection is to prevent runaway trains and/or wagons to pass into a route. This is achieved by setting nearby turnouts to a position that points “away” from the route. Example: PIC The upper turnout, of course, needs to be locked in straight (normal) position, while the lower one is not relevant for the route itself. But what if the lower turnout was set to the diverging (reverse) position and the driver of another train approaching signal B fails to see the red light? This train would crash into the first one. To minimise danger, that other train would need to be routed towards signal D. There are, of course, situations, where both positions of a turnout would conflict with a route equally. In those situations, there’s nothing you can do and no flank lock needs to be set.

3.4 Assigning main signals to TCBs

Main signals in the advtrains interlocking system are positioned - like in real life - at the border of track sections, because routes also start and end there. For advtrains to know from which signal which routes can be set, you need to assign the signal to a TCB. To do this, perform the following steps:

If not already happened, set up a TCB (you don’t need to, but are advised to, configure track sections there)

Place the signal a few meters in front of the TCB, so that trains stopping at the signal do never pass the TCB

Locate the side of the TCB which points in the direction that trains will proceed past the signal, as shown in the figure below.

Right-click the TCB, and click “Assign a signal” on this side.

Punch the signal.

PIC If you haven’t set an influence point for the signal yet, the influence point formspec automatically opens. You can assign a signal to each side of a TCB. This is, for example, useful when creating block sections on a bi-directional main running line.

Only main signals can ever be assigned to TCBs, because static ones can either not display “Danger” or do not permit to proceed at all.

3.5 Shunt

routes

The information in this section is subject to future change because of safety issues! Operating railways is not all about driving trains around. Coupling, decoupling and moving single engines, wagons or groups of wagons across a station, called shunting, also plays an important role. Remember what we said about routes: There must be no rail vehicles on the route. So what if you have some goods wagons ready on a siding, and want to couple an engine to it? You can not set a regular route into the siding, because it is occupied. The solution is to program a second route into the siding, but with the difference that it already ends at the rear-facing signal of it, so it doesn’t include the siding section itself: PIC The Sht2 route then needs to show a shunt aspect, which instructs the driver to proceed slowly and watch out for vehicles on the route. To show a “free” aspect here would be wrong, because that would mean that the track is free until the next main signal, which it is clearly not. Note that advtrains_interlocking currently does not allow to set individual aspects for routes, this is a feature still to be implemented soon. Shunt routes like this are, so far, the only exception to the “Routes should end at a signal facing the same direction” rule.

3.6 Route Release

In early real-life interlocking systems, routes either had to be cancelled by the signalman after the train had passed the route, or there was a single release contact at the end of the route. However, as interlocking systems evolved and the position of trains is now roughly known by the track sections, portions of the route can be freed as soon as the train has left the corresponding section. AdvTrains has chosen a modern approach to route releasing. Each turnout lock is associated to a track section belonging to the route’s path. Once the train leaves this section, all assigned locks are also freed. Please note that reversing a train outside of stations is not only discouraged, but also very dangerous, because even real-world interlocking system do not expect this. There is a clear, human-sense rule that you should never reverse the driving direction of a train while on a main line or on a turnout. Else, you can be considered a terrorist. (quote from professional!)

3.7 Programming a route

The route programming procedure is quite straightforward if you’ve read the previous sections and understood how routes should be set. Routes always start at a main signal. You must have assigned the signal to a TCB, as described earlier. When you right-click the main signal, it no longer changes its aspect. Instead, a formspec pops up, showing you an (empty) list of routes with the possibility to set them or to create new routes. Click the “Create new route” button to start programming a new route. The form closes, and an arrow is displayed on the TCB. You are now in “Route Programming” mode, programming the first track section of the route. Now:

Put any turnouts you need to lock in the correct position (e.g. by right-clicking them). This includes flank protection. Punch them. This makes a marker with a blue lock symbol appear. If you punch a turnout again, or punch the marker, you can remove the lock again. When you’ve locked all turnouts in the current section, go to and punch the TCB that is the border to the next track section the train proceeds into.

Depending on the situation, you are now offered some possibilities to proceed:

Click the “Advance to next section” button if your route consists of more sections with turnouts to lock, and you need to continue programming. Follow the above steps to set locks for the next section.

Once you’ve clicked the “Advance” button, the lock markers change to a red lock symbol, telling they can’t be changed anymore. Repeat the above procedure until you are ready to complete the programming procedure:

Click the “Finish route HERE” button when you’ve set up the locks for the last track section of the route and punched the final TCB (the one with the next signal). You will be asked for a route name and your route will be saved. The “Finish route at end of NEXT section” button (third button) is an useful quickhand to make the route proceed one more section. Using this button is equivalent to first clicking the “Advance” button, then flying to the end of the next track section and finishing the route there. You can not (officially) set turnout locks in the final section using this method.

A few hints:

If one turnout should be locked by more than one section, set the lock only in the last of those sections. Locking the same turnout in multiple sections of a single route results in undefined behavior! If you accidentally advanced the route wrongly, you can use the “Step back one section” button to undo this. If you want to stop programming the entire route without saving it, use the “Cancel route programming” button. The third button is especially useful for programming simple block sections on a main running line, since you can stay at the starting signal (punch starting TCB and select third button). If a route should end in a dead end, you MUST use the “Finish in NEXT section” button, because there is no final TCB that you could punch. The third button does NOT work on sections with more than 2 exits, because the system won’t be able to determine the final TCB of the route then.

4 Interlocking system operation

Setting up the interlocking for a portion of a railway network requires some time, experience and planning, but once done, there’s not much to do anymore to make trains run on your, now safer, railway. This section covers some useful practices to route trains across your network.

At the moment, routes can either be set by clicking the signal or via LuaATC, or by using the “Remote Routesetting” button from the Onboard Computer. It is planned to control this via a “signal box” view based on the currently broken itrainmap.

4.1 Train Safety System

The Train Safety System, called “LZB” in the code (from the german term Linienzugbeeinflussung, although this is a completely different system), ensures that trains obey any restrictions imposed by signals when influence points are set. This way, it is not possible to pass signals at danger or to bypass speed restrictions. It is possible to overrun red signals, if a route is cancelled while a train is approaching. Real interlocking systems use a mechanism called Approach locking for this, however, as of now, there’s no similar system in this mod. If a red signal is overrun, the train brakes using emergency brake (“BB”) and can not be moved any further. You should then examine the situation and drive the train backwards out of the section. As of now, changing the driving direction of a train always clears any imposed speed restrictions.

4.2 Simple

route setting and cancelling

To set a route, simply right-click the signal, select a route and click “set route”. If there are no conflicts, the signal turns green and the train is allowed to proceed. It may be possible that the route can not be set, because one or more other routes conflict with the current one, or a section is blocked. In this case, the signal stays red, and the conflicting item is shown in the formspec. As soon as the conflict is resolved (by cancellation or release of the conflicting route, or the section becoming free), the requested route will be set and the signal turns green. If a route is either requested or set, it can be cancelled from the signalling formspec. This means that all turnouts and sections are released, and the signal reverts back to red. This of course only works when the train has not passed the signal yet. There is no mechanism for Approach Locking.

4.3 Automatic Working

Block signals on main running lines usually only have a single route to set, the one proceeding along the main line. Their purpose is only to show whether there are trains in the next section. So, it would be convenient if this only route would set itself again after a train passed. This is what Automatic Working is for. Set a route, click “Enable Automatic Working”, and as soon as a train passes, the route is automatically re-set. This function is nearly identical to SimSig automatic signals. It can also be useful on a line with high traffic, when there’s a low-frequented access to a siding. You’d enable automatic working for the main route and cancel it only when you need a train to go into the siding.

5 Final notes

The interlocking system is mainly finished, though there are still some plans and ideas. They include:

Signalbox panels, as revival of itrainmap

Individual signal aspects for routes

Distant signals

On-Train head-up display for oncoming signals (they have something like this in Czech Republic, I forgot how it’s called.) Apart from this, there’s the large oncoming project of a new timetable-based train automation system, but this will take some time to evolve and is out of the scope of this document. If you have any suggestions, corrections, improvements, criticism or cute kittens and stuff, you can always contact me by various means (Forum PM, E-Mail (orwell@bleipb.de), Linuxworks server chat a.s.o.). Have fun!

- orwell